OpenStudy (anonymous):

Calculus 2 Help-A rancher has property along a river that he needs to divide into three separate, equal-sized regions: one for the cows, one for the horses, and one for the sheep. The rancher has 1000 ft of available fencing with which to divide up the land. There will need to be enough fencing for three outer sides of the property as well as two lengths of fencing to divide the region into three parts. Determine the values of x and y, and the area of an individual pen so that its area is maximized. Be sure to use calculus to justify you have found a maximum.

OpenStudy (anonymous):

So i got that the equation will be F=1000ft=4x+y and A=1/3xy

OpenStudy (amistre64):

Is the river included as one side of the pens? cause in alot of problems it is..

OpenStudy (anonymous):

you dont need fence for the river part

OpenStudy (anonymous):

so no

OpenStudy (amistre64):

then we need 4x + y = 1000, right?

OpenStudy (anonymous):

yea the 2 equations i set up are my up top

OpenStudy (amistre64):

good, but I need to work it thru in my head too:)

OpenStudy (anonymous):

no prob

OpenStudy (anonymous):

take your time

OpenStudy (amistre64):

and the area of each pen is x(y/3)

OpenStudy (amistre64):

Since y = 1000 - 4x, can we use that in the area?

OpenStudy (anonymous):

yea perfect

OpenStudy (amistre64):

the deriviate and solve for max right?

OpenStudy (amistre64):

esxucse typos, my keyboard hates me:)

OpenStudy (anonymous):

no prob, yea

OpenStudy (amistre64):

i get (3000 - 24x)/9 for a derivative...what about you?

OpenStudy (anonymous):

hold on, did you set it up first by substituting?

OpenStudy (amistre64):

x = 3000/24 = 1500/12 = 750/6..... yeah substitute in y=1000-4x.

OpenStudy (anonymous):

so A(x)=1000x-4X^2 all over 3

OpenStudy (amistre64):

x = 375/3...x = 125 for me

OpenStudy (anonymous):

sorry im going slow

OpenStudy (amistre64):

its ok...gives me time for my mind to wander :)

OpenStudy (anonymous):

was my A(x) correct?

OpenStudy (amistre64):

Your A(x) was correct...

OpenStudy (anonymous):

then you took the derivative?

OpenStudy (amistre64):

yep, the derivative tells us where we can find bends inthe graph that are our max and min values.

OpenStudy (amistre64):

You remember your quotient rule?

OpenStudy (anonymous):

yea let me do it one sec

OpenStudy (anonymous):

3000-24x all over 9

OpenStudy (amistre64):

yep; and theres another way we could have approached it. 1/3 is a constant so we can pull it out and just derivative (1000x-4x^2). Either way is the same

OpenStudy (anonymous):

y did i get x=27000/24

OpenStudy (amistre64):

solve for y' = 0. What makes (3000-24x)/9 = 0?

OpenStudy (amistre64):

the top has to equal 0 for ay of it to matter; so throw away the 9

OpenStudy (amistre64):

3000-24x = 0 what is x?

OpenStudy (anonymous):

so 1000-8x

OpenStudy (amistre64):

thats good too

OpenStudy (amistre64):

1000 - 8x = 0...is the same thing

OpenStudy (anonymous):

1000/8=x

OpenStudy (anonymous):

125=x

OpenStudy (amistre64):

yep, now reduce it or use it as is if you want :) Good job

OpenStudy (anonymous):

y=1000-4(125)

OpenStudy (amistre64):

yep....

OpenStudy (anonymous):

A=1/3(500)*125

OpenStudy (amistre64):

(500)(125)/3 is more accurate, but yeah...

OpenStudy (anonymous):

20833ft^2, that seems large?

OpenStudy (amistre64):

when talking about football field measurements, I imagine it would be :)

OpenStudy (anonymous):

oh ok , now the other part is the maximum

OpenStudy (anonymous):

2nd derivative right?

OpenStudy (amistre64):

total inclosed area is 62500 ft^2

OpenStudy (amistre64):

20833 1/3 each

OpenStudy (anonymous):

they just want individual

OpenStudy (anonymous):

so 20833ft^2 per pen

OpenStudy (amistre64):

we already found the maximum... there is only one hump in a quadratic..

OpenStudy (amistre64):

you can second derive it if you want, just to prove it to yourself, but it aint needed

OpenStudy (anonymous):

i need to do all steps she takes off for everything

OpenStudy (anonymous):

Is this 100% right, she had like inequality equations and domain stuff on her examples. I am confused

OpenStudy (amistre64):

lol..when y'' is (+) we have a concave down which is an upside down bowl on the graph. And that is the max.

OpenStudy (amistre64):

its right....so far :) use the math luke.....the maaatthhhh...

OpenStudy (anonymous):

k haha

OpenStudy (amistre64):

might wanna double check that concavity thing, I never use it enough to remember if Im right :)

OpenStudy (anonymous):

soundds good

OpenStudy (anonymous):

it should be negative to be a max <0

OpenStudy (amistre64):

remember CAVE....CAVE UP is (+) and CAVE DOWN is (-) or the other way around..

OpenStudy (anonymous):

i mean in concavity relation

OpenStudy (anonymous):

if its >0, its concave up oh okay

OpenStudy (amistre64):

lol....ill trust you on that :)

OpenStudy (anonymous):

how do you have a max then

OpenStudy (amistre64):

picture the graph of (-x^2) thats all this is.

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

hold up is the second derivative -8

OpenStudy (amistre64):

thats right, so the cave is down. its an upside down bowl.

OpenStudy (amistre64):

-(x^2) ... y'=-(2x)....y'' = -(2)

OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):

sorry, lol i am so bad at optimization

OpenStudy (amistre64):

:)

OpenStudy (anonymous):

how do the domains work

OpenStudy (anonymous):

she wants us to show that too D:(0, )

OpenStudy (amistre64):

there is nothing limiting our domain, so it canbe any real number (R)

OpenStudy (amistre64):

look at the fractions we made, if anything makes the bottom a 0, throw it out. Nothing matters, so keep it all

OpenStudy (anonymous):

so (0, Infinity)

OpenStudy (amistre64):

I suppose we should focus on positive numbers to make sense of it all, so yeah. except it should be [0,inf) 0 is included.

OpenStudy (amistre64):

wait....just had athought

OpenStudy (amistre64):

we only have so much fencing to work with... so make the D [0,1000]

OpenStudy (amistre64):

or you might wanna think about like this: we need 4 amounts of x so ...maybe...[0,1000/4] ? take a gander at it and see :)

OpenStudy (anonymous):

she likes simplicity

OpenStudy (amistre64):

my gut tells me that we do [0,1000]..

OpenStudy (anonymous):

my friend said the domain is 0 to 250

OpenStudy (amistre64):

does your friend tell you how to find the max area :)

OpenStudy (anonymous):

its your x

OpenStudy (anonymous):

He said solve 1000x-4x^2 for x

OpenStudy (anonymous):

and thats your A( )>0

OpenStudy (anonymous):

which is your max by the Extreme value theorem

OpenStudy (amistre64):

thats fine.....well use 0 to 250......its as good as anyother I can think of :)

OpenStudy (anonymous):

no he said solve the inequality 1000-4x>0

OpenStudy (anonymous):

thats what the 250 is

OpenStudy (amistre64):

1000/4 = 250....i know iknow...yeahyeahyeah....

OpenStudy (anonymous):

can we walk through it step by step, I am so confused

OpenStudy (anonymous):

is 125 okay because it falls in the domain?